Two new parents decide that they will need to save $375,000 to pay for their son's college education. They start saving as soon as he is born. If they invest in a mutual fund that pays about 9.2% interest per year, how much will they need to put away each month to have enough by the child's 18th birthday?
Answer #1

Using financial calculator, Set BGN=1
Input: FV = 375000

I/Y=9.2/12

N=18*12=216

Solve for PMT as -678.45

Monthly amount to be saved at the start of each month = $678.45